OpenStudy (anonymous):

An equilateral triangle has an altitude of 15 m. What is the perimeter of the triangle?

OpenStudy (anonymous):

just the answer or explanation?

OpenStudy (anonymous):

P=30sqrt(3)

OpenStudy (anonymous):

a = 15 * 2/ sqrt(3) = 17.3

OpenStudy (anonymous):

=> P = 3 * 17.3 = 51.96